Calculate the pH of a solution formed by mixing 65 mL of 0.19 M NaHCO3 with 75 mL of 0.17 M Na2CO3. Express your answer using two decimal places.
Answer #1

total volume = 65 + 75 = 140 ml = 0.14 L

concentration of NaHCO3 = 0.19 x 65 / 140 = 0.088 M

concentration of Na2CO3 = 0.17 x 75 / 140 = 0.091M

now mixture of NaHCO3 and Na2CO3 act as acidic buffer.

for acidic buffer

pH = pKa + log [Na2CO3] / [NaHCO3]

pKa of NaHCO3 = 10.3

pH = 10.3 + log[0.091] / [0.088]

pH = 10.31